You are managing a portfolio of $1 million. Your target duration is 10 years, and you can invest in two bonds, a zero-coupon bond with maturity of five years and a perpetuity, each currently yielding 10.0%.

a. What weight of each bond will you hold to immunize your portfolio? (Round your answers to 2 decimal places.)

b. How will these weights change next year if target duration is now nine years? (Round your answers to 2 decimal places.)
Answer #1

a.

Duration of Zero Coupon Bond = 5 years

Duration of Perpetuity = 1.10/0.10 = 11 years

Let weight of zero coupon bond be w,

So,

10 = 5w + (1 - w)(11)

w = 0.1667

Weight of Zero Coupon Bond = 16.67%

Weight of Perpetuity = 1 - 0.1667 = 83.33%

b.

Duration of Zero Coupon Bond = 4 years

Duration of Perpetuity = 1.10/0.10 = 11 years

Let weight of zero coupon bond be w,

So,

9 = 4w + (1 - w)(11)

w = 0.2857

Weight of Zero Coupon Bond = 28.57%

Weight of Perpetuity = 1 - 0.2857 = 71.43%